OpenStudy (anonymous):

Can someone check my answer if I got it right. For f(x)=-4x^2 + 6x - 1 and g(x)=7x-3. Find the following:

jimthompson5910 (jim_thompson5910):

that is incorrect

jimthompson5910 (jim_thompson5910):

\[\Large (f-g)(x)=f(x)-g(x)\] \[\Large (f-g)(x)=(-4x^2 + 6x - 1)-(7x-3)\] \[\Large (f-g)(x)=-4x^2 + 6x - 1-7x+3\] I'll let you finish up.

OpenStudy (anonymous):

so it -4x^2 -x+2

jimthompson5910 (jim_thompson5910):

correct